# Bayesian Structure Learning in Undirected Gaussian Graphical Models: Literature Review with Empirical Comparison

### v2

@@ 3.1.3 MCMC-based approaches on the precision matrix space
Methods on the space of precision matrices remove the necessity of sampling a graph at every iteration. They create an MCMC chain $(\mathbf{K}^{(1)},\cdots,\mathbf{K}^{(S)})$ solely over the space of positive definite $p\times p$ matrices. The methods are computationally efficient and have appealing theoretical properties. We also refer to these types of methods as K-methods. K-methods do not use a prior on the graph and are, therefore, not able to incorporate a prior belief in the graphical structure. Moreover, since K-methods do not provide MCMC samples in the graph space, they still require some operation to select a graph $G$ from the posterior samples of the precision matrix. Two examples of such operations are thresholding and credible intervals. The thresholding method (Wang, 2012) first computes the edge inclusion probability $p_{e}$ of an edge $e$ and then includes it in the estimate of the graph if $p_{e}>\gamma$ where $\gamma\in(0,1)$ is some threshold. The credible intervals method (Li et al., 2019a) first constructs intervals, for every edge $(i,j)$, that contains $\gamma\%$ of the values $\{k_{ij}^{(1)},k_{ij}^{(2)},\dots,k_{ij}^{(S)}\}$ for some user-defined $\gamma$. An edge is then excluded if and only if the interval includes zero.
@@
For K-methods, the prior on the precision matrix $P(\mathbf{K}|G)$ no longer depends on $G$ and can therefore be written as $P(\mathbf{K})$. We call these priors K-only priors. Most K-only priors are defined by two probability distributions. One for off-diagonal elements and one for diagonal elements. The elements of K-only priors are not independent due to the positive definite constraint and these priors also contain an intractable normalizing constant. Specific choices of K-only priors enable algorithms to be computationally efficient and/or have attractive theoretical properties. All K-methods are listed in Table 3 with their corresponding priors.
@@
All K-methods share one or more of three useful theoretical properties. First, all K-methods meet the balance condition and therefore converge to the posterior distribution of the precision matrix (Tierney, 1994, Theorem 1). Second, most K-methods have a connection with the frequentist graphical lasso algorithm. Specifically, priors of K-methods are often chosen such that the precision matrix that maximizes the posterior $P(\mathbf{K}|\mathbf{X})$ is equal to the graphical lasso estimate obtained by solving the optimization problem (5) for some regularization parameter $\lambda$ and penalty function $p(\mathbf{K})$. The third and last theoretical property only holds for the K-Horseshoelike and the K-Low rank and diagonal (LRD) methods and builds off a theorem originally introduced by Banerjee & Ghosal (2015). It states that the posterior probability of precision matrices $\mathbf{K}$ that are “far away” from the true precision matrix $\mathbf{K}^{*}$ goes to zero, as the number of observations goes to infinity. Or, put differently, the posterior distribution $P(\mathbf{K}|\mathbf{X})$ is concentrated around the true value $K^{*}$ (Sagar et al., 2024, Theorem 4.6). Table 3 contains all K-methods with their corresponding theoretical properties.
Reference Name Value of $P(k_{ij})$ Theor. Prop.
@@
Sagar et al. (2024) K - Horseshoe-like Horseshoe-like $\propto 1$
@@
Table 3: K-methods for Bayesian structure learning with their corresponding reference, name, K-only prior, and theoretical properties. Note that all marginal K-only priors are not equal to the distributions mentioned in the table due to the positive definite constraint. The horseshoe and horseshoe-like distribution are defined in the corresponding articles. The theoretical properties are i) the Markov chain converges to the posterior distribution, ii) the maximum a posteriori estimate equals the maximum likelihood estimate (5), and iii) the posterior concentrates around the true value. Papers with a $*{}^{*}$ are not yet published.
@@ 3.2 Expectation Conditional Maximization Approaches
MCMC-based approaches are popular due to the wide variety of inferences they can provide using Equation (14). They are, however, computationally demanding because of the large parameter space they explore. Expectation Conditional Maximization (ECM) methods are a more efficient alternative. They allow for the quick computation of the maximum a posteriori (MAP) estimate of the precision matrix but do not provide inference on the rest of the posterior. The three ECM methods in Bayesian structure learning are listed in Table 4.
Every iteration $s=1,\dots,S$ of an ECM algorithm renders a new precision matrix $\mathbf{K}^{(s)}$. After $S$ iterations, where $S$ is large enough, the sequence $(\mathbf{K}^{(1)},\mathbf{K}^{(2)},\cdots,\mathbf{K}^{(S)})$ converges to the MAP estimator. In each iteration two steps are applied: the Expectation step (E-step) and the Conditional-Maximization step (CM-step). In iteration $s$, the E-step evaluates $Q(\mathbf{K}|\mathbf{K}^{(s-1)})$, defined as the expected value of the log posterior with respect to the current conditional distribution of $G$ given $\mathbf{K}$ and $\mathbf{X}$. In the CM-step the new precision matrix $\mathbf{K}^{(s)}$ is set to the matrix that optimizes $Q(\mathbf{K}|\mathbf{K}^{(s-1)})$. The pseudo-code is shown in Algorithm 4.
@@
Output: A sequence $\mathbf{K}^{(0)},\mathbf{K}^{(1)},..,\mathbf{K}^{(S)}$ that converges to the MAP estimate.
Algorithm 4 Expectation Conditional Maximization algorithm
ECM methods in Bayesian structure learning differ in three aspects: their prior, their theoretical properties, and their output. See Table 4 for an overview. First, ECM methods define their prior on the precision matrix $P(\mathbf{K}|G)$ using three distributions. One for elements $k_{ij}$ for which $(i,j)\in G$, one for elements $k_{ij}$ for which $(i,j)\not\in G$, and one for diagonal elements. Like the G-Wishart prior, SS prior, and K-only priors discussed before, the elements $k_{ij}$ are not independent, because the matrix $\mathbf{K}$ needs to be positive definite. The priors therefore also contain an intractable normalizing constant. Second, ECM methods provide consistent results on the graph and/or the precision matrix. For example, with their ECM - Bayesian regularization for Graphical models with Unequal Shrinkage (ECM-BAGUS) method Gan et al. (2019) show that the estimate of the precision matrix $\hat{\mathbf{K}}$ gets arbitrarily close to the true precision matrix $\mathbf{K}^{*}$ as the number of observations goes to infinity. To be specific, under certain assumptions and for a constant $C$, they show that
@@
where $||\cdot||_{\infty}$ denotes the $l_{\infty}$ norm. The ECM-Horseshoelike (ECM-HL) method by Sagar et al. (2024) proves a similar result. The ECM - BAGUS method is the only ECM method that also provides a consistent result for the graph. Third, like the K-methods of Section 3.1.3, ECM methods require some post-hoc operation to provide inference on the graph. The ECM-BAGUS method is thereby the only ECM method that outputs the edge inclusion probabilities (13).
A limitation of ECM methods is that they only work for specifically designed priors on the precision matrix. Sagar et al. (2023) present a method that can output a MAP estimate of the precision matrix for any prior on the precision matrix, as long as the prior densities of the off-diagonal elements are completely monotone. Completely monotone priors form a large class of priors that include the priors of the K-Horseshoe and ECM-HL methods.
Reference Name Value of $P(k_{ij}|G)$ Theory Output
@@
Table 4: ECM methods for Bayesian structure learning with their corresponding reference, name, prior on $\mathbf{K}$, theoretical properties, and output. DE$(a,b)$ denotes the double exponential distribution with mean $a$ and scale parameter $b$. N$(a,b)$ denotes the normal distribution with mean $a$ and standard deviation $b$. Exp$(\eta)$ is the exponential distribution with rate $\eta$. Note that all marginal priors are not equal to the distributions mentioned in the table due to the positive definite constraint. The theoretical properties are i)the MAP estimate is consistent, and ii) the estimate of the graph is consistent. The output is the MAP estimate of the precision matrix ($\mathbf{K}_{MAP}$) and/or the edge inclusion probabilities ($p_{e}$) of (13).
@@ 3.3 Regression Approaches
In Section 2.2, we saw that structure learning is equivalent to finding the coefficients of $p$ different regressions. That is, estimating the coefficients $\beta_{ij}$ in (3) allows for direct estimation of the graphical model, since $\beta_{ij}=\frac{k_{ij}}{k_{ii}}$ and hence $\beta_{ij}=0\iff(i,j)\not\in G$. Regression methods use a Bayesian approach to find the coefficients $\beta_{ij}$. Regression methods offer a simple approach to Bayesian structure learning. The resulting estimates of the graph and precision matrix, however, come without any theoretical guarantees.
There are two regression methods: The R - Bayesian Lasso Neighborhood Regression Estimate (R - BLNRE) method by Li & Zhang (2017) and the R - Projection (R-P) method by Williams et al. (2018). Both put priors on each coefficient $\beta_{ij}$ and create Markov chains $(\beta_{ij}^{(1)},\cdots,\beta_{ij}^{(S)})$ for each $\beta_{ij}$. These MCMC samples are approximately samples from the posterior distribution $P(\beta_{ij}|\mathbf{X})$. Like MCMC-based approaches on the space of precision matrices (Section 3.1.3), regression methods now use a post-MCMC operation to obtain an estimate of the graph and precision matrix.
The R-BLNRE method finds the $\beta_{ij}$ in (3) using the Bayesian lasso, a popular Bayesian algorithm for finding regressions coefficients by Park & Casella (2008). In the Bayesian lasso the prior on each coefficient $\beta_{ij}$ is set to the double exponential distribution. The mode of the resulting posterior distribution $P(\beta_{ij}|\mathbf{X})$ equals the frequentist estimate obtained by solving the optimization problems in (4).
The R-P method by Williams et al. (2018) uses the horseshoe prior for the regression coefficients. The same prior is used for the elements of the precision matrix in the K- Horseshoe method by Li et al. (2019a). Using so-called projection predictive selection, the samples from the resulting posterior $P(\beta_{i}|\mathbf{X})$ are used to output an estimate of the graph and precision matrix.
@@ 3.4 Hypothesis Approaches
Hypothesis methods calculate the edge inclusion probabilities in (13) directly. They thereby circumvent the time-consuming exploration of the model space necessary for MCMC-based methods. To do so, they first formulate, for every $(i,j)$, a null hypothesis $H_{0}:k_{ij}=0$ and an alternative hypothesis $H_{1}:k_{ij}\not=0$. Then, they calculate the Bayes factor in favour of $H_{1}$
$BF_{ij}$ $=\frac{P(\mathbf{X}|H_{1})}{P(\mathbf{X}|H_{0})}$
[…]
This strategy was introduced by Giudici (1995), who derived a closed-form expression of $BF_{ij}$ when the number of observations is bigger than the number of variables, i.e., $p<n$. The Hypothesis Leday and Richardson (H - LR) method by Leday & Richardson (2018) builds upon this result by deriving a closed-form expression for $BF_{ij}$ that also holds for $p>n$ and show that it is consistent, i.e., $\lim_{n\to\infty}BF_{ij}=0$, if $H_{0}$ is true and $\lim_{n\to\infty}BF_{ij}=\infty$, if $H_{1}$ is true. This Bayes factor, however, is not scale-invariant. To correct for this, the H-LR method first scales $BF_{ij}$ to obtain the scaled Bayes factor $sBF_{ij}$. Let $P(sBF_{ij}>b|H_{0})$ denote the tail probability that one observes a scaled Bayes factor higher than some constant $b$, given that $H_{0}$ is true. When $H_{0}$ is true, $sBF_{ij}$ follows a beta distribution. The tail distribution $P(sBF_{ij}>b|H_{0})$ can thus be computed exactly. The edge inclusion probabilities now follow by setting $P((i,j)\in G)=1-P(sBF_{ij}>b|H_{0})$. For a prior on the precision matrix they use the Wishart distribution, which is equivalent to the G-Wishart distribution $P(\mathbf{K}|G)$ in Equation (10) when $G$ is the complete graph. The H-LR method does not use a prior on the graph and is therefore, like the K-methods, not able to incorporate a prior belief on the graphical structure.
The Hypothesis - BGGM (H-BGGM) method by Williams & Mulder (2020) formulates the hypothesis using the partial correlations $\rho_{ij}$ in (1). It is named after their R package Bayesian Gaussian Graphical Models (BGGM). The resulting hypotheses are $H_{0}:\rho_{ij}=0$, $H_{1}:\rho_{ij}>0$, and $H_{2}:\rho_{ij}<0$. They also formulate an unrestricted hypothesis $H_{u}:\rho_{ij}\in(-1,1)$. These hypotheses allow for the computation of a scale-invariant Bayes factor. The edge exclusion probability $P((i,j)\not\in G)$ can now be calculated as follows:
@@
Here, $BF_{ku}=\frac{P(\mathbf{X}|H_{k})}{P(\mathbf{X}|H_{u})}=\frac{P(H_{k}|\mathbf{X})}{P(H_{k})}$ for $k=0,1,2$. $P(H_{k})$ is the prior distribution for accepting hypothesis $H_{k}$. The Bayes factor $BF_{ku}$ is therefore just a ratio of the posterior and prior of the hypotheses $H_{k}$. The authors compute this ratio using three priors. First, using the Wishart prior, they show that if the precision matrix $K$ has a Wishart prior, the prior and posterior partial correlations approximately follow normal distributions. The Bayes factors $BF_{ku}$ can then be computed in two ways: analytically or by sampling repeatedly from the Wishart distribution and using Equation 1. The Wishart distribution, however, is not able to incorporate prior beliefs on the graph. Therefore, they propose two other types of prior distributions for the precision matrix: the Corrected Wishart (CW) prior and the matrix-F prior. For both prior distributions, they show that the Bayes factor $BF_{ku}$ can be computed using the so-called Savage-Dickey ratio by Dickey (1971). The H-BGGM method can also be used for confirmatory hypothesis tests. These kinds of hypothesis tests involve more than one partial correlation; for example, $H_{0}:\rho_{12}>\rho_{34}$ and $H_{1}:\rho_{12}\leq\rho_{34}$.

### v3

@@ 3.1.3 MCMC-based approaches on the precision matrix space
Methods on the space of precision matrices remove the necessity of sampling a graph at every iteration. They create an MCMC chain $(\mathbf{K}^{(1)},\dots,\mathbf{K}^{(S)})$ solely over the space of positive definite $p\times p$ matrices. The methods are computationally efficient and have appealing theoretical properties. We also refer to these types of methods as K-methods. K-methods do not use a prior on the graph and are, therefore, not able to incorporate a prior belief in the graphical structure. Moreover, since K-methods do not provide MCMC samples in the graph space, they still require some operation to select a graph $G$ from the posterior samples of the precision matrix. Two examples of such operations are thresholding and credible intervals. The thresholding method (Wang, 2012) first computes the edge inclusion probability $p_{e}$ of an edge $e$ and then includes it in the estimate of the graph if $p_{e}>\gamma$ where $\gamma\in(0,1)$ is some threshold. The credible intervals method (Li et al., 2019a) first constructs intervals, for every edge $(i,j)$, that contains $\gamma\%$ of the values $\{k_{ij}^{(1)},k_{ij}^{(2)},\dots,k_{ij}^{(S)}\}$ for some user-defined $\gamma$. An edge is then excluded if and only if the interval includes zero.
@@
For K-methods, the prior on the precision matrix $P(\mathbf{K}|G)$ no longer depends on $G$ and can therefore be written as $P(\mathbf{K})$. We call these priors K-only priors. Most K-only priors are defined by two probability distributions. One for off-diagonal elements and one for diagonal elements. The elements of K-only priors are not independent due to the positive definite constraint and these priors also contain an intractable normalizing constant. Specific choices of K-only priors enable algorithms to be computationally efficient and/or have attractive theoretical properties.
@@
All K-methods share one or more of three useful theoretical properties. First, all K-methods meet the balance condition and therefore converge to the posterior distribution of the precision matrix (Tierney, 1994, Theorem 1). Second, most K-methods have a connection with the frequentist graphical lasso algorithm. Specifically, priors of K-methods are often chosen such that the precision matrix that maximizes the posterior $P(\mathbf{K}|\mathbf{X})$ is equal to the graphical lasso estimate obtained by solving the optimization problem (5) for some regularization parameter $\lambda$ and penalty function $p(\mathbf{K})$. The third and last theoretical property is the concentration of the precision matrix around the true value as in the SS-BC algorithm by Jalali et al. (2020) (see also Section 3.1.1). Table 3 contains all K-methods with their corresponding K-only priors and theoretical properties.
Reference Name $P(k_{ij})$ Theor. Prop.
@@
Sagar et al. (2024) K - Horseshoe-like Horseshoe-like $\propto 1$
Table 3: K-methods for Bayesian structure learning with their corresponding reference, name, K-only prior, and theoretical properties. Note that all marginal K-only priors are not equal to the distributions mentioned in the table due to the positive definite constraint. The horseshoe and horseshoe-like distribution are defined in the corresponding articles. The theoretical properties are i) the Markov chain converges to the posterior distribution, ii) the maximum a posteriori estimate equals the maximum likelihood estimate (5), and iii) the posterior concentrates around the true value. Papers with a are not yet published.
K-methods are not scale-invariant, i.e. scaling the observations of one or more variables affects the posterior distribution. K-methods typically deal with this by standardizing the data to have unit variance. Carter et al. (2023) observe, however, that this standardization can adversely affect inference. Instead, Carter et al. (2023) propose to perform inference directly on the partial correlations and define a class of priors for which the resulting posterior is scale-invariant.
@@ 3.2 Expectation Conditional Maximization Approaches
MCMC-based approaches are popular due to the wide variety of inferences they can provide using Equation 14. They are, however, computationally demanding because of the large parameter space they explore. Expectation Conditional Maximization (ECM) methods are a more efficient alternative. They allow for the quick computation of the maximum a posteriori (MAP) estimate of the precision matrix but do not provide inference on the rest of the posterior. The three ECM methods in Bayesian structure learning are listed in Table 4.
At every iteration $s=1,\dots,S$, an ECM algorithm renders a new precision matrix $\mathbf{K}^{(s)}$. The sequence $(\mathbf{K}^{(1)},\mathbf{K}^{(2)},\dots)$ converges to the MAP estimator. In each iteration two steps are applied: the Expectation step (E-step) and the Conditional-Maximization step (CM-step). In iteration $s$, the E-step evaluates $Q(\mathbf{K}|\mathbf{K}^{(s-1)})$, defined as the expected value of the log posterior with respect to the current conditional distribution of $G$ given $\mathbf{K}$ and $\mathbf{X}$. In the CM-step the new precision matrix $\mathbf{K}^{(s)}$ is set to the matrix that optimizes $Q(\mathbf{K}|\mathbf{K}^{(s-1)})$. The pseudo-code is shown in Algorithm 4.
@@
Output: A sequence $\mathbf{K}^{(0)},\mathbf{K}^{(1)},\dots,\mathbf{K}^{(S)}$ that converges to the MAP estimate.
Algorithm 4 Expectation Conditional Maximization (ECM) algorithm
ECM methods in Bayesian structure learning differ in three aspects from each other: their prior, their theoretical properties, and their output. See Table 4 for an overview. First, each ECM method uses a different SS prior (see Subsection 3.1.1). Second, each ECM method provides consistent results on the graph and/or the precision matrix. For example, with their ECM-BAGUS method Gan et al. (2019) show that the estimate of the precision matrix $\hat{\mathbf{K}}$ gets arbitrarily close to the true precision matrix $\mathbf{K}^{*}$ as the number of observations goes to infinity. To be specific, under certain assumptions and for a constant $C$, they show that
@@
where $||\cdot||_{\infty}$ denotes the $l_{\infty}$ norm. The ECM-Horseshoelike (ECM-HL) method by Sagar et al. (2024) proves a similar result. The ECM - BAGUS method is the only ECM method that also provides a consistent result for the graph. Third, ECM methods differ in the output they provide. Like the K-methods of Section 3.1.3, ECM methods require some post-hoc operation to provide inference on the graph. The ECM-BAGUS method is thereby the only ECM method that outputs the edge inclusion probabilities (13).
A limitation of ECM methods is that they only work for specifically designed priors on the precision matrix. This is overcome by Sagar et al. (2023), who present a method that outputs a MAP estimate of the precision matrix for a large class of priors.
Reference Name $P(k_{ij}|G)$ Theory Output
@@
Table 4: ECM methods for Bayesian structure learning with their corresponding reference, name, prior on $\mathbf{K}$, theoretical properties, and output. DE$(a,b)$ denotes the double exponential distribution with mean $a$ and scale parameter $b$. N$(a,b)$ denotes the normal distribution with mean $a$ and standard deviation $b$. Exp$(\eta)$ is the exponential distribution with rate $\eta$. Note that all marginal priors are not equal to the distributions mentioned in the table due to the positive definite constraint. The theoretical properties are i) the MAP estimate is consistent, and ii) the estimate of the graph is consistent. The output is the MAP estimate of the precision matrix ($\mathbf{K}_{MAP}$) and/or the edge inclusion probabilities $p_{e}$ of (13).
@@ 3.3 Regression Approaches
In Section 2.2, we saw that structure learning is equivalent to finding the coefficients of $p$ different regressions. That is, estimating the coefficients $\beta_{ij}$ in (3) allows for direct estimation of the graphical model, since $\beta_{ij}=\frac{k_{ij}}{k_{ii}}$ and hence $\beta_{ij}=0\iff(i,j)\not\in G$. Regression methods use a Bayesian approach to find the coefficients $\beta_{ij}$, thereby offering a simple approach to Bayesian structure learning. The resulting estimates of the graph and the precision matrix, however, come with limited theoretical guarantees.
There are two regression methods: The R - Bayesian Lasso Neighborhood Regression Estimate (R - BLNRE) method by Li & Zhang (2017) and the R - Projection (R-P) method by Williams et al. (2018). Both put priors on each coefficient $\beta_{ij}$ and create Markov chains $(\beta_{ij}^{(1)},\dots,\beta_{ij}^{(S)})$ for each $\beta_{ij}$. These MCMC chains provide approximate samples from the posterior distribution $P(\beta_{ij}|\mathbf{X})$. Like MCMC-based approaches on the space of precision matrices (Section 3.1.3), regression methods now use a post-MCMC operation to obtain an estimate of the graph and the precision matrix.
The R-BLNRE method finds the $\beta_{ij}$ in (3) using the Bayesian lasso, a popular algorithm by Park & Casella (2008). In the Bayesian lasso the prior on each coefficient $\beta_{ij}$ is set to the double exponential distribution. The mode of the resulting posterior distribution $P(\beta_{ij}|\mathbf{X})$ equals the frequentist estimate obtained by solving the optimization problems in (4).
The R-P method by Williams et al. (2018) uses the horseshoe prior for the regression coefficients. The same prior is used for the elements of the precision matrix in the K- Horseshoe method by Li et al. (2019a). Using so-called projection predictive selection, the samples from the resulting posterior $P(\beta_{ij}|\mathbf{X})$ are used to output an estimate of the graph and precision matrix.
@@ 3.4 Hypothesis Approaches
Hypothesis methods circumvent the time-consuming exploration of the model space necessary for MCMC-based methods. Instead, they perform inference on the graph directly. To do so, they first formulate, for every $(i,j)$, a null hypothesis $H_{0}:k_{ij}=0$ and an alternative hypothesis $H_{1}:k_{ij}\not=0$. Then, they calculate the Bayes factor in favour of $H_{1}$, given by
$BF_{ij}$ $=\frac{P(\mathbf{X}|H_{1})}{P(\mathbf{X}|H_{0})}.$
This strategy was introduced by Giudici (1995), who derived a closed-form expression of $BF_{ij}$ when the number of observations is larger than the number of variables, i.e., $n>p$. The Hypothesis Leday and Richardson (H-LR) method by Leday & Richardson (2018) builds upon this result by deriving a closed-form expression for $BF_{ij}$ that also holds for $p>n$ and show that it is consistent, i.e., $\lim_{n\to\infty}BF_{ij}=0$ if $H_{0}$ is true, and $\lim_{n\to\infty}BF_{ij}=\infty$ if $H_{1}$ is true. The H-LR method scales $BF_{ij}$ to obtain a scale-invariant Bayes factor $sBF_{ij}$. By computing $sBF_{ij}$ for all edges and selecting a threshold $\gamma$, one can obtain a graph estimate $\hat{G}$ by including an edge $(i,j)$ in $\hat{G}$ if $sBF_{ij}>\gamma$. For a prior on the precision matrix, they use the Wishart distribution, which is equivalent to the $G$-Wishart distribution $P(\mathbf{K}|G)$ in Equation 10 when $G$ is a complete graph. The H-LR method does not use a prior on the graph and is therefore, like the K-methods, not able to incorporate a prior belief on the graphical structure.
The Hypothesis - BGGM (H-BGGM) method by Williams & Mulder (2020) formulates the hypothesis using the partial correlations $\rho_{ij}$ in (1). The resulting hypotheses are $H_{0}:\rho_{ij}=0$, $H_{1}:\rho_{ij}>0$, and $H_{2}:\rho_{ij}<0$. They also formulate an unrestricted hypothesis $H_{u}:\rho_{ij}\in(-1,1)$. These hypotheses allow for the computation of a scale-invariant Bayes factor. The edge exclusion probability $P((i,j)\not\in G)$ can now be calculated as follows:
@@
Here, $BF_{ku}=\frac{P(\mathbf{X}|H_{k})}{P(\mathbf{X}|H_{u})}=\frac{P(H_{k}|\mathbf{X})}{P(H_{k})}$ for $k=0,1,2$. $P(H_{k})$ is the prior distribution for accepting hypothesis $H_{k}$. The Bayes factor $BF_{ku}$ is therefore just a ratio of the posterior and prior of the hypotheses $H_{k}$. Using the Wishart prior for the precision matrix, they show that the prior and posterior partial correlations approximately follow normal distributions. The Bayes factors $BF_{ku}$ can then be computed analytically or by sampling repeatedly from the Wishart distribution and using Equation 1. The Wishart distribution, however, is not able to incorporate prior beliefs on the graph. Therefore, they propose two other priors: the Corrected Wishart prior and the matrix-F prior. The H-BGGM method can also be used for confirmatory hypothesis tests. These kinds of hypothesis tests involve more than one partial correlation; for example, $H_{0}:\rho_{12}>\rho_{34}$ and $H_{1}:\rho_{12}\leq\rho_{34}$.